Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: In B configuration; able to support 16 baseline board pairs (for up to 16 subbands). New software over the last week (Executor, MCAF); currently an issue with archive extraction - to be resolved ASAP.
  • Key dates:
    • 07-10 Mar: New Worlds, New Horizons (Santa Fe)
    • 07 May - 15 Jun: 4-band science observing
  • Staff Travel:
  • Visitors:
    • 05 Jan - 06 Apr: John Cannon sabbatical at DSOC
    • 22 Feb - 08 Mar: Sonja Vrcic (Correlator)
    • 28 Feb - 28 Mar: Bob Sault

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • EA13 returns to array (24Mar; EA18 moves out).
    • L band on EA16 (Wednesday)
    • Issues: Baseline determination still showing some large offsets, weather suspected. Previous issue resolved:
The recent move into B array uncovered some oddities in the  fitting 
of antenna positions ('baselines') in the AIPS routine LOCIT. This was
traced to the format of the AIPS OTT= 'Over the Top' table written by
BDF2AIPS - it was in real*4 instead of real*8. This has been fixed by
Eric Greisen.

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon
    • All day: Science Operations/testing observations
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0630: Science operations
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations
  • Thu, Fri:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: fshift sub-band anti-aliasing causing phantom spectral lines when using recirculation for narrow bands (only worry if Michael has to prep your file); Work this week with Bob et al.
  • Correlator: Missing/mislabeled scans in data sets (small percentage but recurring).
  • Correlator: time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags);fix available pending testing verification.
  • Array/Construction: Dhawan/Rupen: RF switch issues; new set coming to be tested.
  • Correlator: Delay rocking across bands. Brent investigating; tests to come.
  • Correlator: Subset of BlBs misbehaving (only affects hardware stacking modes). Currently restricted to 16 working Baseline boards; several RSRO projects are pending.
  • Antenna/High Freq: Reference pointing accuracy (C, X, Ku). Rick has noted that we are worse than before - larger fraction of observations indicate larger pointing offsets (Perley).
  • Array: Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).

Commissioning Groups

polarizations, spanning the entire range between 1 and 50 GHz (for 3C48, 3C138, 3C147, 3C286).

Testing This week

  • Array: Sowinski: Ku-band collimation
  • Array: Dhawan: Baseline determination run
  • Correlator: Rupen: Subarray testing
  • Correlator: Rupen: Scan time alignment testing (dumptrig epoch)
  • Correlator: Sault/Clark/Rupen: fshift/spectral splatter testing
  • Solar: Perley/Bastian: Solar observing tests (20dB testing)
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Spectral line: Momjian: Zeeman test follow-up


RSRO B configuration status:

Individual reports (updates)


Demo Science Status

Her A Status report (Perley)



Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised. - Sault visit * Low Priority [TBD]: Phase (and amp) wobbles; long history.fshift investigation.
-- JosephMcMullin - 2011-02-28
Topic revision: r3 - 2011-03-02, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback